# Massless higher-spin supermultiplets in 5D harmonic superspace

### v1

@@ Appendix C Prepotentials for 5D conformal supergravity
Prepotentials for the 4D conformal supergravity in the harmonic superspace framework were studied in Ref. GIOS ; Galperin:1987ek ; KT ; Butter:2010sc ; Butter:2015nza . In this Appendix, we uplift this construction to the 5D harmonic superspace and introduce the prepotentials describing 5D Weyl supermultiplet, although the action for 5D conformal supergravity is not known.
@@ C.1 Conformal transformations in harmonic superspace
Conformal transformations in the 5D harmonic superspace were studied in Ref. Kuzenko2006 . In this subsection, we collect basic relations for the superconformal Killing vector and the corresponding transformations of superfields.
@@
By definition, an iso-tensor superfield $H^{i_{1}\ldots i_{n}}=H^{(i_{1}\ldots i_{n})}$ is said superconformal primary of weight $p/2$ if it transforms by the rule
@@
By definition, superconformal transformation of a primary analytic superfield $\Phi^{(q)}$ with conformal weight $n$ and $U(1)$ charge $q$ is
@@ C.2 Superconformal properties of the free $q$-hypermultiplet model
In the previous subsection, we defined rigid superconformal transformations in 5D harmonic superspace. In this section, we begin by considering transformations of the hypermultiplet superfield under local superconformal group. Invariance of the free hypermultiplet model under rigid superconformal symmetry will be discussed at the end of this subsection.
@@
The general solution to these constraints is expressed via single unconstrained gauge parameter $l^{--}$:
@@
In general, this variation is non-vanishing, unless we impose additional constraint on the operator $\Lambda$:
@@
Below, we will show that the invariance of the $q$-hypermultiplet model under the local superconformal transformations (338) is achieved through coupling to conformal supergravity background.
@@ C.3 Prepotentials for conformal supergravity
By definition, acting on an analytic superfield $\Phi_{A}$, this operator should preserve analyticity, $D^{+}_{\hat{\alpha}}(H^{++}\Phi_{A})=0$. As a corollary, the prepotentials $h^{++M}$ obey the constraints
@@
$\lambda$-gauge transformations of the supergravity prepotentials are derived from supergravity gauge transformation of the covariant harmonic derivative, $\delta{\cal D}^{++}=[{\cal D}^{++},\Lambda]-\lambda^{++}D^{0}$:
@@
Scalar superfield prepotential $\mathfrak{U}$ with gauge freedom of the form (364) was introduced in Ref. Butter:2014xxa for 5D conformal supergravity.
If one chooses to work with the gauge condition (356), the residual gauge freedom make it possible to bring the Zupnik prepotential $\Psi^{\hat{\alpha}-}$ to the form
@@ C.4 Hypermultiplet coupling and supercurrent
Under the hypermultiplet transformations (338) the actions (42) and (366) vary as follows
@@
Variation of this action over $U$ yields the conserved supercurrent in the free hypermultiplet model:

### v2

@@ Appendix C Prepotentials for 5D conformal supergravity
Prepotentials for 4D conformal supergravity in the harmonic superspace framework were studied in Ref. GIOS ; Galperin:1987ek ; KT ; Butter:2010sc ; Butter:2015nza . In this Appendix, we uplift this construction to the 5D harmonic superspace and introduce the prepotentials describing 5D Weyl supermultiplet, although the action for 5D conformal supergravity is not known.
@@ C.1 Conformal transformations in harmonic superspace
Conformal transformations in the 5D harmonic superspace were studied in Ref. Kuzenko2006 . In this subsection, we collect basic relations for the superconformal Killing vector and the corresponding transformations of the superfields.
@@
By definition, an iso-tensor superfield $H^{i_{1}\ldots i_{n}}=H^{(i_{1}\ldots i_{n})}$ is said to be superconformal primary of weight $p/2$ if it transforms by the rule
@@
By definition, a superconformal transformation of a primary analytic superfield $\Phi^{(q)}$ with conformal weight $n$ and $U(1)$ charge $q$ is
@@ C.2 Superconformal properties of the free $q$-hypermultiplet model
In the previous subsection, we defined rigid superconformal transformations in 5D harmonic superspace. In this section, we begin by considering transformations of the hypermultiplet superfield under the local superconformal group. Invariance of the free hypermultiplet model under rigid superconformal symmetry will be discussed at the end of this subsection.
@@
The general solution to these constraints is expressed via a single unconstrained gauge parameter $l^{--}$:
@@
In general, this variation is non-vanishing, unless we impose an additional constraint on the operator $\Lambda$:
@@
Below, we will show that the invariance of the $q$-hypermultiplet model under the local superconformal transformations (354) is achieved through coupling to the conformal supergravity background.
@@ C.3 Prepotentials for conformal supergravity
By definition, when acting on an analytic superfield $\Phi_{A}$, this operator should preserve analyticity, $D^{+}_{\hat{\alpha}}(H^{++}\Phi_{A})=0$. As a corollary, the prepotentials $h^{++M}$ obey the constraints
@@
$\lambda$-gauge transformations of the supergravity prepotentials are derived from a supergravity gauge transformation of the covariant harmonic derivative, $\delta{\cal D}^{++}=[{\cal D}^{++},\Lambda]-\lambda^{++}D^{0}$:
@@
The scalar superfield prepotential $\mathfrak{U}$ with gauge freedom of the form (380) was introduced in Ref. Butter:2014xxa for 5D conformal supergravity.
If one chooses to work with the gauge condition (372), the residual gauge freedom makes it possible to bring Zupnik’s prepotential $\Psi^{\hat{\alpha}-}$ to the form
@@ C.4 Hypermultiplet coupling and supercurrent
Under the hypermultiplet transformations (354), the actions (47) and (382) vary as follows
@@
Variation of this action with respect to $U$ yields the conserved supercurrent in the free hypermultiplet model: